# Characterization of the directed landscape from the KPZ fixed point

### v2

@@ Abstract
We show that the directed landscape is the unique coupling of the KPZ fixed point from all initial conditions at all times satisfying three natural properties: independent increments, monotonicity, and shift commutativity. Equivalently, we show that the directed landscape is the unique directed metric on $\mathbb{R}^{2}$ with independent increments and KPZ fixed point marginals. This gives a framework for proving convergence to the directed landscape given convergence to the KPZ fixed point. We apply this framework to prove convergence to the directed landscape for a range of models, some without exact solvability: asymmetric exclusion processes with potentially non-nearest neighbour interactions, exotic couplings of ASEP, the random walk and Brownian web distance, and directed polymer models. All of our convergence theorems are new except for colored ASEP and the KPZ equation, where we provide alternative proofs.
@@ 1 Introduction
The KPZ (Kardar-Parisi-Zhang) universality class is a loose term for a collection of $1$-dimensional models of random interface growth and $2$-dimensional random metric or polymer models that exhibit the same behaviour under rescaling. The past twenty-five years have seen a period of intense and fruitful research on this class, with progress propelled by the discovery of a handful of exactly solvable models, including exclusion processes, directed polymers, last passage percolation models, and the KPZ equation itself. See [Qua12, Cor12, Rom15, BG16, Zyg22, Gan22] and references therein for background on the KPZ universality class and related areas.
The KPZ fixed point is the universal scaling limit for growth models in the KPZ class, and the directed landscape is the universal scaling limit for metrics and polymer models. The KPZ fixed point can be constructed as a marginal of the directed landscape, and one may optimistically hope the opposite is also true: that we can construct the directed landscape from the KPZ fixed point. This is the goal of the present paper.
@@
To describe the KPZ fixed point, we first consider one particular family of growth models: exclusion processes on $\mathbb{Z}$. Let $p$ be a measure on $\mathbb{Z}$ with finite support. An exclusion process on $\mathbb{Z}$ with jump distribution $p$ is an interacting particle system where particles occupy sites in $\mathbb{Z}$, with at most one particle per site. A particle at site $x$ will attempt to jump to site $x+v$ at rate $p(v)$, succeeding if the site $x+v$ is unoccupied (the exclusion rule). All particle jumps are independent, and to avoid degeneracies we assume that the symmetrized support $\{v\in\mathbb{Z}:p(v)+p(-v)>0\}$ additively generates $\mathbb{Z}$. We can associate to an exclusion process at time $t$ a height function $h_{t}:\mathbb{Z}\to\mathbb{Z}$, by letting
$h_{t}(x+1)=\begin{cases}h_{t}(x)+1,\quad\text{there is a particle at site $x$},\\ h_{t}(x)-1,\quad\text{there is a hole at site $x$}.\end{cases}$
This rule determines the height function up to an anchor point at $h_{t}(0)$. While we allow ourselves flexibility in choosing $h_{0}(0)$, in order to define the height functions consistently as $t$ changes, we stipulate that $h_{t}(0)=h_{t-}(0)-2$ if a particle jumps from the interval $(-\infty,-1]$ to the interval $[0,\infty)$ at time $t$, and that $h_{t}(0)=h_{t-}(0)+2$ if a particle jumps from $[0,\infty)$ to $(-\infty,-1]$ at time $t$. Otherwise we set $h_{t}(0)=h_{t-}(0)$.
When the expectation $\sum_{v\in\mathbb{Z}}vp(v)\neq 0$, we call such a process an asymmetric exclusion process (AEP). Without loss of generality, we may assume this expectation is positive, and by rescaling time, that it equals $1$. In an AEP, the height function decreases at a linear rate, and so the process $(h_{t})_{t\geq 0}$ produces an growing interface where the growth is governed by independent forcing, but is constrained by the local structure of the interface through the exclusion rule. These features are typical of growth models in the KPZ universality class. The special case when the jump distribution $p$ is supported on $\{-1,1\}$ is known as the asymmetric simple exclusion process (ASEP). If $p(1)=1$ and $p(v)=0$ for all $v\neq 1$ then this is the totally asymmetric simple exclusion process (TASEP). TASEP height function evolutions are particularly easy to understand: the height function evolves by a flipping local maxima to a local minima, $h_{t}(x)\mapsto h_{t}(x)-2$, at rate $1$.
Matetski, Quastel, and Remenik [MQR21] showed that for TASEP, the process
$h^{\epsilon}_{t}(x):=\epsilon^{1/2}h_{2\epsilon^{-3/2}t}(2\epsilon^{-1}x)+\epsilon^{-1}t$
converges as $\epsilon\to 0$ to a limiting Markov process whose state space is the set of upper semicontinuous functions from ${}\mathbb{R}\to\mathbb{R}\cup\{-\infty\}$ satisfying certain growth conditions. This limiting process is the KPZ fixed point. Convergence to the KPZ fixed point has since been verified for a handful of models, including solvable last passage percolation models [NQR20, DV21b], the KPZ equation and polymer models [QS23, Vir20], the AEPs described above [QS23], and more recently the stochastic six-vertex model [ACH24a].
@@
The KPZ universality class also contains $2$-dimensional models of random metrics naturally built from a family of independent random variables or a noise on the plane. These models can be connected back to models of random growth in various ways, for example by considering the process of distances from a point $o$ to a collection of lines indexed by time $t$. Varying $o$ gives rise to a coupling of many different random growth processes. Conversely, if we are given a particularly natural coupling of random growth processes simultaneously from all initial conditions, then we can often realize this coupling through a random metric.
To make this discussion more concrete, consider Liggett’s basic coupling for exclusion processes started from different initial configurations [Lig76]. In the basic coupling for AEPs with jump distribution $p$, we can view each pair of sites $n,n+v\in\mathbb{Z}$ as being attached to an independent rate-$p(v)$ Poisson clock $\Pi_{n,v}$. When the Poisson clock $\Pi_{n,v}$ rings, in all of the AEPs, if there is a particle at site $n$ then this particle will attempt to jump to site $n+v$, succeeding whenever the exclusion rule allows it. We can also couple AEPs started at different times with this coupling. This coupling is equivalent to colored or multi-type AEP, see [AAV11, ACH24b] for background.
In the special case of TASEP, this coupling can be fully realized with a directed metric construction on $\mathbb{Z}\times\mathbb{R}$. For each $x\in\mathbb{Z}$, let $\Delta_{x}:y\mapsto-|x-y|$ be the narrow wedge centred at $x\in\mathbb{Z}$. Define the function
@@
for any $s\leq t$ and $x,y\in\mathbb{Z}$. Here the right hand side is the height function at time $t$ and location $y$ for TASEP started from $\Delta_{x}$ at time $s$ in the basic coupling through the Poisson clocks $\Pi=\{\Pi_{n,v}\}$. Set $d_{\Pi}(x,s;y,t)=-\infty$ for $t>s$. We have $d_{\Pi}(p;p)=0$ for all $p\in\mathbb{Z}\times\mathbb{R}$ and the function $d_{\Pi}$ satisfies a reverse triangle inequality
@@
This makes $d_{\Pi}$ a directed metric, see Section 2 for background. Moreover, for any initial condition $h_{0}$, the TASEP evolution of $h_{0}$ started at time $s$ in the coupling governed by the clocks $\Pi$ is given by the variational formula
@@
Note that if we replace TASEP by a general ASEP in (1) then formula (3) no longer holds, and if we further replace ASEP with a general AEP, then we lose the triangle inequality (2) as well.
@@
then $d_{\Pi}^{\epsilon}$ converges as $\epsilon\to 0$ to the random function $\mathcal{L}:\mathbb{R}^{4}_{\uparrow}\to\mathbb{R},\mathbb{R}^{4}_{\uparrow}:=\{(x,s;y,t)\in\mathbb{R}^{4}:s<t\}$ known as the directed landscape, constructed in [DOV22] two years after the construction of the KPZ fixed point and with a completely different method. For this particular coupling of TASEP, this convergence was recently proven by Aggarwal, Corwin, and Hegde [ACH24b]. The directed landscape has also been shown to be the scaling limit of classical last passage percolation models [DOV22, DV21b], the KPZ equation [Wu23], and the stochastic six-vertex model [ACH24b]. The directed landscape is connected to the KPZ fixed point through a limiting version of (3). For an upper semi-continuous function $f:\mathbb{R}\to\mathbb{R}\cup\{-\infty\}$, the process
@@
is a KPZ fixed point started from $f$ at time 00, and so the directed landscape gives a coupling of the KPZ fixed point, simultaneously from all times and all initial conditions.
@@ 1.1 Main results: characterization theorems
    The choice of $\mathbb{Q}$ is arbitrary, both in the above theorem and throughout the paper. We could replace it with any countable dense subset of $\mathbb{R}$. Note also that we could have replaced $\mathbb{Q}^{4}_{\uparrow}$ with $\mathbb{Q}^{4}_{\uparrow}\cap(\mathbb{R}\times I)^{2}$ for any open interval $I\subset\mathbb{R}$; the proof does not change and we would recover a directed landscape defined only on the strip ×$\mathbb{R}\times I$.
@@
  3. 3.
    The proof of Theorem 1.1 does not actually use much precise information about the directed landscape $\mathcal{L}$. Indeed, all we really need is that $\mathcal{L}$ satisfies the three conditions in Theorem 1.1 along with the following stationarity property:
    Property $\ast$:
We also give an alternate version of Theorem 1.1 that characterizes the directed landscape as the unique coupling of the KPZ fixed point satisfying certain natural conditions. This version, while not difficult to show given Theorem 1.1, may better demonstrate how the characterization links back to couplings of discrete models such as the AEP basic coupling described above.
Here and throughout this paper, we use $\operatorname{UC}$ to denote the space of all upper semicontinuous $f:\mathbb{R}\to\bar{\mathbb{R}}:=\mathbb{R}\cup\{\pm\infty\}$, and let $\operatorname{UC}_{0}\subset\operatorname{UC}$ be the subspace of all $f$, satisfying $f(x)<\infty$ for any $x\in\mathbb{R}$, and:
$\limsup_{x\to\pm\infty}\frac{f(x)}{x^{2}}\leq 0.$ (5)
@@
    (Independent increments) For any collection of disjoint intervals $[s_{i},t_{i}],i=1,\dots,k$, the operators $\mathcal{K}_{s_{i},t_{i}},i=1,\dots,k$ are independent.
@@
    To check that assumptions $2-4$ above are natural, we can compare with properties in the basic coupling of AEPs. The independent increment property in this setting is immediate. Shift commutativity is also immediate, and comes from the flexibility we initially had in defining our height function at 00 given the particle description. Deterministic monotonicity can be readily checked for ASEP, but is not true for general AEPs. However, an approximate version still holds and passes to monotonicity in the limit (as we will show in Section 8.3).
@@
    Max-preservation fails for many prelimiting models (e.g. even for ASEP, AEP in the basic coupling), and we do not have a simple argument for how it arises in the limit. Rather, we prove this through utilizing the full power of Theorem 1.1 and the variational characterization (4) of the law of the KPZ fixed point.
In the statement of Theorem 1.3, we have restricted our initial conditions to the (essentially arbitrary) dense set $\operatorname{NW}$ mostly for convenience. Other initial conditions come along for the ride, as exhibited by the following corollary.
@@
Note that condition (7) will always hold with $\mathbb{Q}$ replaced by some countable dense set $Q$ which may depend on $f$. For example, letting $Q=\{\operatorname*{arg\,max}_{x\in[q,q^{\prime}]}f(x):q<q^{\prime}\in\mathbb{Q}\}$ gives such a set, where here we choose the rightmost $\operatorname*{arg\,max}$ if multiple points exist. Therefore in practice Corollary 1.5 can be applied to any function in $\operatorname{UC}_{0}$.
@@ 1.2 Main results: new framework for directed landscape convergence
The practical upshot of Theorems 1.1 and 1.3 is that if we can show that a model converges to the KPZ fixed point, then we can typically show that it converges to the directed landscape by checking a few straightforward conditions. This offers a new approach to proving convergence to the directed landscape. Note that all previous approaches to studying convergence rely on extra combinatorial structure coming from either some generalization of the Robinson-Schensted correspondence [DOV22, DZ24, DV21b, Wu23, DV24], or the Yang-Baxter equation [ACH24b].
@@
The following models converge to the directed landscape after appropriate rescaling: the O’Connell-Yor polymer (Theorem 7.4), the KPZ equation (Theorem 7.2), the random walk web distance and Brownian web distance as conjectured in [VV23] (Theorem 6.5), AEPs under the basic coupling (Theorem 8.7), and ASEPs and TASEPs under various exotic couplings (Theorem 8.4).
For the ASEP exotic couplings (to be defined in Definition 8.2), we note that the particle coupling, where Poisson clocks are attached to labelled particles rather than sites, is included in this collection.
@@
    The only models above where directed landscape convergence has been previously established are the KPZ equation [Wu23], and colored ASEP/TASEP [ACH24b] (which is a special case of both our AEP convergence theorem and our exotic coupling theorem for ASEP).
    For the many other models covered in Theorem 1.6, including the random walk/Brownian web distance, general AEPs under the basic coupling, and ASEP under general exotic couplings, such directed landscape convergence seems beyond the scope of all previous approaches, since the necessary precise combinatorial structures (such as line ensembles or the Yang-Baxter equation) are unknown or absent.
@@
    For brevity, we do not attempt to be comprehensive in applying our convergence framework. Beyond the models in Theorem 1.6, there are various other models where our convergence framework should apply straightforwardly. For example, for integrable last passage percolation (LPP) models, where directed landscape convergence was previously shown in [DOV22, DV21b], our Theorems 1.1 and 1.3 could be applied to get alternative proofs, using a priori known KPZ fixed point convergence (e.g. [MQR21] or [Liu22] for exponential LPP, [NQR20] for Brownian LPP, [JR21] for geometric LPP). Our framework could also be applied to obtain directed landscape convergence for variants of TASEP where the original method of [MQR21] has been extended to find formulas suitable for proving convergence to the KPZ fixed point, see [Ara20, MR23a, MR23b].

### v3

@@ Abstract
We show that the directed landscape is the unique coupling of the KPZ fixed point from all initial conditions satisfying three natural properties: independent increments, monotonicity, and shift commutativity. Equivalently, we show that the directed landscape is the unique directed metric on $\mathbb{R}^{2}$ with independent increments and KPZ fixed point marginals. This unifies the two central objects in the KPZ universality class.
Our main theorem also provides a general framework for proving convergence to the directed landscape given convergence to the KPZ fixed point. We apply this framework to prove landscape convergence in a range of models: exotic couplings of ASEP and TASEP, the random walk and Brownian web distances, and a class of non-integrable asymmetric exclusion processes with the basic coupling that perturb off of TASEP (this final class requires random initial data). All of our convergence theorems are new except for colored TASEP, where we provide a short alternative proof.
@@ 1 Introduction
The KPZ (Kardar-Parisi-Zhang) universality class is a broad family of one-dimensional random growth models and two-dimensional random metric or polymer models that exhibit the same behaviour under rescaling. The past twenty-five years have seen a period of intense and fruitful research on this class, propelled by the discovery of a handful of exactly solvable models, including certain exclusion processes, directed polymers, last passage percolation, and the KPZ equation itself. For background, we refer the reader to [Qua12, Cor12, Rom15, BG16, Zyg22, Gan22] and references therein.
The KPZ fixed point is the universal scaling limit for growth models in the KPZ class, and the directed landscape is the universal scaling limit for metrics and polymer models. The KPZ fixed point can be constructed as a marginal of the directed landscape, and one may optimistically hope the reverse is also true: that the directed landscape can be constructed from the KPZ fixed point. This is the goal of the present paper.
@@
To describe the KPZ fixed point, we first consider one particular growth model: the asymmetric simple exclusion process (ASEP). ASEP is an interacting particle system on $\mathbb{Z}$, but for our purposes here it will be more convenient to define it directly as a height function evolution (see Section 7 for the connection to the particle system). Define
$\operatorname{SRW}=\{h:\mathbb{Z}\to\mathbb{Z}:h(i)-h(i-1)=\pm 1\text{ for all }i\}.$
This is the space of two-sided simple random walk paths on $\mathbb{Z}$. ASEP is a Markov process $(h_{t},t\geq 0)$ on $\operatorname{SRW}$ where $h_{t^{-}}(x)\mapsto h_{t}(x)+2$ at rate $p$ if $x\in\mathbb{Z}$ is a local maximum of $h_{t^{-}}$ and $h_{t^{-}}(x)\mapsto h_{t}(x)-2$ at rate $p+1$ if $x\in\mathbb{Z}$ is a local minimum of $h_{t^{-}}$ (here $h_{t^{-}}=\lim_{s\to t^{-}}h_{s}$). This is a natural interface growth model, where there is an asymmetry (depending on $p$) causing the interface to drift downwards at constant rate over time. We have normalized our rates the downward drift is speed $1/2$ for any $p$.
The special case when $p=0$ is referred to as TASEP, the totally asymmetric simple exclusion process. TASEP is one of the most tractable models in the KPZ universality class, and Matetski, Quastel, and Remenik [MQR21] showed that
$h^{\epsilon}_{t}(x):=\epsilon^{1/2}h_{2\epsilon^{-3/2}t}(2\epsilon^{-1}x)+\epsilon^{-1}t$ (1)
converges as $\epsilon\to 0$ to a limiting Markov process $(\mathfrak{h}_{t},t\geq 0)$ known as the KPZ fixed point. The state space for the KPZ fixed point we use in this paper is
$\operatorname{UC}_{0}:=\Big{\{}f:\mathbb{R}\to\mathbb{R}\cup\{-\infty\}:f\text{ upper semicontinuous},f\not\equiv-\infty,\quad\limsup_{x\to\pm\infty}\frac{f(x)}{x^{2}}\leq 0\Big{\}},$ (2)
which is the largest space on which $\mathfrak{h}_{t}$ is almost surely finite at all positive times, see [SV21, Proposition 6.1].
Since the original work of [MQR21], convergence to the KPZ fixed point has been verified for a handful of models, including intregrable last passage percolation (LPP) models [NQR20a, DV21b], the KPZ equation and polymer models [Vir20], certain non-integrable perturbations of TASEP [QS23] (see Theorem 7.6 for a precise statement), and more recently the stochastic six-vertex model and ASEP as described above [ACH24b, ACH24a].
@@
The KPZ universality class also certain $2$-dimensional random metric and polymer models. Marginals of these models define random growth models in the KPZ class, e.g.  we can consider a process of balls of increasing radius started at a fixed point $p$. Varying the base point Conversely, if we are given a particularly natural coupling of random growth processes, then we can often (but not always!) realize this coupling through a random metric or a polymer model.
To make this discussion more concrete, we can consider Liggett’s basic coupling for ASEP [Lig76]. In the basic coupling, each point $x\in\mathbb{Z}$ is associated to two Poisson processes on $\mathbb{R}$: $\Pi_{x,+}$, with intensity $p$, and $\Pi_{x,-}$, with intensity $p+1$. All Poisson processes are independent. Each copy $h_{t}$ of ASEP will attempt the jump $h_{t^{-}}(x)\mapsto h_{t}(x)\pm 2$ at each time in the Poisson process $\Pi_{x,\pm}$, succeeding if $h_{t^{-}}$ has a local minimum/maximum or $x$. In brief, height functions jump together. This coupling is equivalent to colored or multi-type ASEP, see [AAV11, ACH24b] for background.
In the special case of TASEP, this coupling can be fully realized with a directed metric construction on $\mathbb{Z}\times\mathbb{R}$. For each $x\in\mathbb{Z}$, define $\Delta_{x}\in\operatorname{SRW}$ by $\Delta_{x}(y)=-|x-y|$. This is the narrow wedge centred at $x$. Define
@@
for any $s\leq t$ and $x,y\in\mathbb{Z}$. Here the right hand side is the height function at time $t$ and location $y$ for TASEP started from $\Delta_{x}$ at time $s$ in the basic coupling through the Poisson clocks $\Pi=\{\Pi_{x,-}:x\in\mathbb{Z}\}$. Set $d_{\Pi}(x,s;y,t)=$-$\infty$ for $t>s$. We have $d_{\Pi}(p;p)=0$ for all $p\in\mathbb{Z}\times\mathbb{R}$ and the function $d_{\Pi}$ satisfies a reverse triangle inequality
@@
This makes $d_{\Pi}$ a directed metric, see Section 2 for background. Moreover, for any initial condition $h_{0}\in\operatorname{SRW}$, the TASEP evolution of $h_{0}$ started at time $s$ in the basic coupling governed by the clocks $\Pi$ is given by the variational formula
@@
We can actually define the metric $d_{\Pi}$ using a notion of paths and path length. We say that a function $\gamma:[s,t]\to\mathbb{Z}$ of bounded variation is a path from $(x,s)$ to $(y,t)$ if $\gamma(s)=x,\gamma(y)=t$, and we define its length
$|\gamma|_{\Pi}=V(\gamma)+2\#\{r\in(s,t]:r\in\Pi_{\gamma(r),-}\},$
where $V(\gamma)$ is the total variation of $\gamma$. With this definition, $d_{\Pi}(p;q)$ is the maximum length of a path from $p$ to $q$. Almost surely, this maximum always exists, making $d_{\Pi}$ a geodesic space. The metric $d_{\Pi}$ is a prototypical random (directed) metric in the KPZ class: a planar metric built from paths defined through a field of i.i.d. random variables or a noise on the plane.
If we replace TASEP by a general ASEP in (3) then the triangle inequality (4) is still valid. However, the metric $d_{\Pi}$ is no longer a geodesic space, and the variational formula (5) (which is closely related to the geodesic property) no longer holds.
@@
then $d_{\Pi}^{\epsilon}$ converges as $\epsilon\to 0$ to the random function $\mathcal{L}:\mathbb{R}^{4}_{\uparrow}\to\mathbb{R},\mathbb{R}^{4}_{\uparrow}:=\{(x,s;y,t)\in\mathbb{R}^{4}:s<t\}$ known as the directed landscape, constructed in [DOV22] two years after the construction of the KPZ fixed point and with a completely different method. For the basic coupling of TASEP and ASEP, directed landscape convergence was recently proven by Aggarwal, Corwin, and Hegde [ACH24b]. Since [DOV22], convergence to the directed landscape has also been shown for classical last passage percolation models [DOV22, DV21b], the KPZ equation [Wu23], and the stochastic six-vertex model [ACH24b]. The directed landscape is connected to the KPZ fixed point through a limiting version of (5). For $f\in\operatorname{UC}_{0}$, the process
@@
is a KPZ fixed point started from $f$ at time 0, and so the directed landscape gives a coupling of the KPZ fixed point, simultaneously from all times and all initial conditions.
@@ 1.1 Main results: characterization theorems
    The choice of $\mathbb{Q}$ is arbitrary, both in the above theorem and throughout the paper. We could replace it with any countable dense subset of $\mathbb{R}$. Note also that we could have replaced $\mathbb{Q}^{4}_{\uparrow}$ with $\mathbb{Q}^{4}_{\uparrow}\cap(\mathbb{R}\times I)^{2}$ for any open interval $I\subset\mathbb{R}$; the proof does not change and we would recover a directed landscape defined only on the strip $\mathbb{R}\times I$.
@@
We also give an alternate version of Theorem 1.1 that characterizes the directed landscape as the unique coupling of the KPZ fixed point satisfying certain natural conditions. This version, while not difficult to show given Theorem 1.1, may better demonstrate how the characterization links back to couplings of discrete models.
@@
    (Independent increments) For any collection of disjoint intervals $s_{i},t_{i}],i=1,\dots,k$, the operators $\mathcal{K}_{s_{i},t_{i}},i=1,\dots,k$ are independent.
@@
    Max-preservation fails for many prelimiting models (e.g. ASEP in the basic coupling), and we do not have a simple argument for how it arises in the limit. Rather, we prove this through utilizing the full power of Theorem 1.1 and the variational characterization (6) of the law of the KPZ fixed point.
    This point is related to the fact that there are many natural prelimiting couplings of random growth models that do not arise from an underlying geodesic space. One upshot of Theorem 5.2 is that geodesics emerge in the limit as long as we started from an underlying coupling with (approximate versions of) the independent increment, monotonicity, and shift commutativity properties.
In the statement of Theorem 1.3, we have restricted our initial conditions to the (essentially arbitrary) dense set $\operatorname{NW}$ mostly for convenience. Other initial conditions come along for the ride.
@@
Note that condition (8) will always hold with $\mathbb{Q}$ replaced by some countable dense set $Q$ which may depend on $f$. For example, letting $Q=\{\operatorname*{arg\,max}_{x\in[q,q^{\prime}]}f(x):q<q^{\prime}\in\mathbb{Q}\}$ gives such a set, where here we choose the rightmost $\operatorname*{arg\,max}$ if multiple points exist. Therefore by Remark 1.4.1, in practice Corollary 1.5 can be applied to any function in $\operatorname{UC}_{0}$.
@@ 1.2 Main results: new framework for directed landscape convergence
A dream of universality.   The central challenge in the field of KPZ—sometimes referred to as the strong KPZ universality problem—is to show that we will see the characteristic 11:22:33 scaling and all KPZ limit objects whenever we start with a microscopic model satisfying a few basic qualitative properties indicating membership in the KPZ class.
At present, we have an excellent understanding of the limit objects in the KPZ class, and there are techniques for showing that models with sufficient exact solvability converge to these limit objects. There are also a few comparison techniques available that allow for small perturbations of exactly solvable models, e.g. see [QS23, Ran23]. However, the problem of analyzing truly non-integrable models (e.g. first and last passage percolation with general weights) seems far out of reach.
Theorem 1.1 can be viewed as providing one small step towards universality. It implies that if we can prove convergence to the KPZ fixed point for a moodel, then we can often upgrade this to the directed landscape convergence by checking a few straightforward conditions. Informally, if you can find a way to get 99% of the way to universality, then Theorem 1.1 can take you the last 1%. Note that all previous approaches to studying landscape convergence rely on extra combinatorial structure coming from either some generalization of the Robinson-Schensted correspondence [DOV22, DZ24b, DV21b, Wu23, DV24], or the Yang-Baxter equation [ACH24b].
To illustrate this idea, we give a handful of simple applications to models where KPZ fixed point convergence has been previously established, but directed landscape convergence has remained elusive.
@@
The following models converge to the directed landscape after appropriate rescaling: the random walk web distance and Brownian web distance as conjectured in [VV23] (Theorem 6.5), ASEPs and TASEPs under various exotic couplings (Theorem 7.4), and general asymmetric exclusion processes in a ‘TASEP-perturbative’ limiting regime and with random initial data (Theorem 7.7).
@@
    If a directed metric model converges to the directed landscape in a sufficiently strong sense, then its geodesics also converge to landscape geodesics, see [DOV22, Section 12] and [DV21b, Section 8]. The required convergence is hypograph convergence. This convergence is weaker than uniform-on-compact convergence and can be established without hard estimates, see [DV21b, Section 10]. We do not explore geodesic or hypograph convergence here, but note that one should be able to conclude geodesic convergence in the above models by applying our approach here with the aformentioned ideas in [DV21b].
@@
    The only two models in Theorem 1.6 where landscape convergence has been previously established are ASEP and TASEP in the basic coupling. For the other models covered in Theorem 1.6, directed landscape convergence seems beyond the scope of all previous approaches, since the necessary combinatorial structures (e.g. line ensembles or the Yang-Baxter equation) are unknown or absent. In particular, the general exclusion processes in Theorem 7.7 gives the first example of a model where there are no exact formulas known for any marginals, but convergence to the directed landscape can be established. The caveat here is that we need to consider exclusion processes with random initial data and in a restrictive (but still highly non-trivial) TASEP-perturbative regime.
  3. 3.
    For brevity, we do not attempt to be comprehensive in applying our convergence framework and have focused on models where no tractable line ensembles are expected to exist. Beyond the models in Theorem 1.6, there are other models where our convergence framework applies straightforwardly. For example, in the original arXiv version of this paper [DZ24a], we included proofs of convergence for the KPZ equation and the O’Connell-Yor polymer starting from the results of [Vir20]. For integrable last passage percolation (LPP) models, where directed landscape convergence was shown in [DOV22, DV21b], our Theorems 1.1 and 1.3 give alternate proofs starting from either KPZ fixed point or multi-time convergence (see [MQR21] or [Liu22] for exponential LPP, [NQR20a] for Brownian LPP, and [JR21] for geometric LPP).
    We expect that our framework can also be applied to obtain directed landscape convergence for variants of TASEP where there are prelimiting formulas that are potentially suitable for taking a KPZ fixed point limit. See [Ara20, NQR20b, MR23a, MR23b] for examples which follow from extensions of the original approach of [MQR21], and [BLSZ23] for inhomogeneous versions of TASEP where an alternate RSK-based approach yields tractable formulas. Note that even pointwise convergence of the kernel has not been shown for these models, other than in a few cases (certain discrete time TASEPs, see [Ara20, MR23a]). In these latter cases, trace class convergence has not been shown.
We finish this section by elaborating on the (T)ASEP exotic couplings in Theorem 1.6. For convenience, we will only couple height function evolutions where $h(0)\in 2\mathbb{Z}$; adding in this condition makes the state space $\operatorname{SRW}$ irreducible under the ASEP evolution. First, we can alternately define the evolution of ASEP with parameter $p$ using two arrays of Poisson processes $Q^{+}(x,h),Q^{-}(x,h)$ on $\mathbb{R}$ indexed by the even lattice
$\mathbb{Z}^{2}_{e}=\{(x,h)\in\mathbb{Z}^{2}:x+h\text{ is even}\}.$
The processes $Q^{+}(x,h)$ have intensity $p$ and the processes $Q^{-}(x,h)$ have intensity $p+1$. An ASEP height function $h_{t}$ will attempt the jump $h_{t^{-}}(x)\mapsto h_{t}(x)\pm 2$ if $t\in Q^{\pm}(x,h_{t^{-}}(x))$ and $h_{t^{-}}$ has a local minimum/maximum at $x$. A pair of arrays $Q^{\pm}$ defines an ASEP evolution when we have the following two conditions:
  • For any $h\in\operatorname{SRW}$ with local minima at $x_{1},\dots,x_{k}$ and local maxima at $y_{1},\dots,y_{\ell}$, the processes
    $Q^{+}(x_{i},h(x_{i})),\;\;i=1,\dots,k,\quad Q^{-}(y_{j},h(y_{j})),\;\;j=1,\dots,\ell$
    are all independent.
  • For any fixed time $t\in\mathbb{R}$, the forward clocks $\{Q^{\pm}(u)\cap[t,\infty):u\in\mathbb{Z}^{2}_{e}\}$ are independent of the backward clocks $\{Q^{\pm}(u)\cap(-\infty,t]:u\in\mathbb{Z}^{2}_{e}\}$.
Now, consider a coupling of ASEP simultaneously from all times and all initial conditions using the same clocks $Q^{-},Q^{+}$. Such a coupling is always monotone, and is jointly Markov by the second bullet point above. In other words, the coupling satisfies prelimiting versions of $2,3$ in Theorem 1.3. There are many ways to impose an additional constraint on the clocks $Q^{\pm}$ so we always satisfy shift commutativity in the limit. To define the exotic couplings in Theorem 1.6, we fix some $u=(a,b)\in\mathbb{Z}^{2}_{e}$ with $b\geq|a|$, and ask that all clocks satisfy the additional condition $Q^{\pm}(u+x)=Q^{\pm}(u)$ for all $x\in\mathbb{Z}^{2}_{e}$. Beyond this, we set the clocks independently. This easily implies shift commutativity in the limit, allowing us to apply Theorem 1.3 to prove directed landscape convergence.
The above construction also shows how our main theorems fail away from the continuum limit. Indeed, there are many distinct exotic couplings of ASEP, which satisfy very precise pre-limiting versions of Theorem 1.3.1-4. All these couplings become equivalent in the KPZ limit. We note also that the exotic coupling construction covers three classic couplings: the case when $u=(0,2)$ is Liggett’s basic coupling, the case when $u=(1,1)$ is the particle coupling, where Poisson clocks are attached to labelled particles rather than sites, and u=(1,1)u=(-1,1) covers the similarly defined hole coupling.
Finally, there are many choices of the clocks $Q^{\pm}$ where shift commutativity does not arise in the limit; here we expect a limiting grand monotone coupling of the KPZ fixed point not driven by the directed landscape. A particularly intriguing example is when all clocks are independent. Here the limiting object should be driven by a three-dimensional, rather than a two-dimensional, noise.
Problem 1.8.
In the above construction, consider the setting where all of the clocks $Q^{\pm}$ are independent. Construct the scaling limit of this model.